# Gorenstein cohomological dimension and stable categories for groups

### v4

@@ 6. Model structure and Stable categories over group rings
We refer to [11] for the notion of exact category, which is commonly attributed to Quillen. By [29, Definition 2.2], an additive category is weakly idempotent complete if every split monomorphism has a cokernel and every split epimorphism has a kernel. The following is immediate.
@@
It is not usually trivial to check that a category has a model category structure, see for example [6, Section 10]. However, thanks to [32, Theorem 2.2] and [29, Theorem 3.3], we have a correspondence between complete cotorsion pairs and the model structure. This brings us a convenience to have the following.
@@
  • the cofibrations (trivial cofibrations) are monomorphisms whose cokernel are cofibrant $RG$-modules (projective $RG$-modules).
@@
For any $M\in\mathcal{F}ib$, we have an exact sequence $0\rightarrow M\rightarrow N\rightarrow L\rightarrow 0$ with $N\in\mathcal{W}$ and $L\in\mathcal{C}of$. Moreover, for $N\in\mathcal{W}$, there is an exact sequence $0\rightarrow K\rightarrow P\rightarrow N\rightarrow 0$, where $P$ is projective and $K\in\mathcal{W}$. Now we consider the pullback of $M\rightarrow N$ and $P\rightarrow N$, and obtain the following commutative diagram.
[…]
We infer $X\in\mathcal{C}of$ from the middle row, where $L\in\mathcal{C}of$ and $P\in\mathcal{P}$. Then, by the left column and the lower row, we infer that the cotorsion pair $(\mathcal{C}of,\mathcal{W})$ is complete.
Consequently, by using [29, Theorem 3.3] we have a model structure on $\mathcal{F}ib$ as stated above, which is corresponding to the triple of classes of $RG$-modules $(\mathcal{C}of,\mathcal{W},\mathcal{F}ib)$. The triple is usually referred to as a Hovey triple, since such a correspondence was obtained by Hovey in [32, Theorem 2.2]. ∎
Let $M$ be an object in a model category. Recall that $M$ is called cofibrant if $0\rightarrow M$ is a cofibration, and it is called fibrant if $M\rightarrow 0$ is a fibration. For the model category $\mathcal{F}ib$, there is a stable category of the Frobenius category $\mathcal{C}of$, whose objects are both cofibrant and fibrant. Moreover, there is an associated homotopy category $\mathrm{Ho}(\mathcal{F}ib)$, which is obtained by formally inverting weak equivalences, that is, the localization of $\mathcal{F}ib$ with respect to the class of weak equivalences.
The following is immediate from a fundamental result about model categories, see for example [31, Theorem 1.2.10].
@@
Note that weak equivalences are crucial in the model category. We have the following characterization, which might be of independent interests.
@@
Let ${\rm StMod}(RG)$ be the stable module category with all fibrant $RG$-modules as objects, and morphisms for any objects $M$ and $N$ are given by complete cohomology of degree zero, that is,
@@
Let $M$ and $N$ be any fibrant $RG$-modules. By the completeness of the cotorsion pair $(\mathcal{C}of,\mathcal{W})$, there exists an exact sequence $0\rightarrow K_{M}\rightarrow Q(M)\rightarrow M\rightarrow 0$, where $Q(M)$ is cofibrant and $K_{M}\in\mathcal{W}$. Hence, the cofibrant approximation $Q(M)\rightarrow M$ is also a trivial fibration, and we refer it (or simply, $Q(M)$) to be a cofibrant replacement of $M$. Then, we have the following isomorphisms
@@
where the first one follows by [31, Theorem 1.2.10(ii)] and Proposition 6.5, and the second one holds by Lemma 6.7.
By basic properties of cofibrant modules (see Proposition 5.5), for fibrant $RG$-modules $M$ and $N$, there exists an integer $r>>0$, such that both $\Omega^{r}(M)$ and $\Omega^{r}(N)$ are cofibrant modules, and moreover, the projective dimension of $K_{M}$ and $K_{N}$ are not more than $r-1$. For $M$, we have exact sequences
$0\longrightarrow\Omega^{r}(M)\longrightarrow P_{r-1}\longrightarrow\cdots\longrightarrow P_{1}\longrightarrow P_{0}\longrightarrow M\longrightarrow 0,$
$0\longrightarrow P^{\prime}_{r}\longrightarrow P^{\prime}_{r-1}\longrightarrow\cdots\longrightarrow P^{\prime}_{1}\longrightarrow Q(M)\longrightarrow M\longrightarrow 0,$
where $P_{i}$ and $P^{\prime}_{i}$ are all projective $RG$-modules; similarly, we obtain such exact sequences for $N$. Moreover, we get the following commutative diagram:
@@
Let ${\rm K}_{ac}(RG\text{-}{\rm Proj})$ and ${\rm K}_{tac}(RG\text{-}{\rm Proj})$ denote the homotopy categories of acyclic complexes and totally acyclic complexes of projective $RG$-modules, respectively. There is a functor $\Omega:{\rm K}_{ac}(RG\text{-}{\rm Proj})\rightarrow\underline{\mathcal{GP}}(RG)$ given by taking the kernel of the boundary map in degree zero.
We summarize some equivalences of triangulated categories as follows, which generalize [34, Theorem 3.10] where the equivalences
@@
Acknowledgements. The author is grateful to Professor I. Emmanouil for sharing the preprint [24] and for helpful comments and suggestions, and also to Professors X.-W. Chen, D.-S. Li and W.-H. Qian for their helpful suggestions. This work is supported by the National Natural Science Foundation of China (No. 11871125).

### v5

@@ 6. Model structure and Stable categories over group rings
We refer to [11] for the notion of exact category, which is commonly attributed to Quillen. By [28, Definition 2.2], an additive category is weakly idempotent complete if every split monomorphism has a cokernel and every split epimorphism has a kernel. The following is immediate.
@@
It is not usually trivial to check that a category has a model category structure, see for example [6, Section 10]. However, thanks to [31, Theorem 2.2] and [28, Theorem 3.3], we have a correspondence between complete cotorsion pairs and the model structure. This brings us a convenience to have the following.
@@
  • the cofibrations (trivial cofibrations) are monomorphisms whose cokernels are cofibrant $RG$-modules (projective $RG$-modules).
@@
Let $M$ be any $RG$-module in $\mathcal{F}ib$. By the above argument we have an exact sequence $0\rightarrow M\rightarrow N\rightarrow L\rightarrow 0$ with $N\in\mathcal{W}$ and $L\in\mathcal{C}of$. Moreover, since ${\rm Gpd}_{RG}M\leq{\rm pd}_{RG}M\otimes_{R}B(G,R)$ is finite, it follows from [29, Theorem 2.10] that there exists an exact sequence $0\rightarrow K\rightarrow H\rightarrow M\rightarrow 0$ in $\mathcal{F}ib$ where $K\in\mathcal{W}$ and $H$ is a Gorenstein projective $RG$-module. By Lemma 5.4, $H$ is also a cofibrant module. Hence, the completeness of the cotorsion pair $(\mathcal{C}of,\mathcal{W})$ follows.
Consequently, by using [28, Theorem 3.3] we have a model structure on $\mathcal{F}ib$ as stated above, which is corresponding to the triple of classes of $RG$-modules $(\mathcal{C}of,\mathcal{W},\mathcal{F}ib)$. The triple is usually referred to as a Hovey triple, since such a correspondence was obtained by Hovey in [31, Theorem 2.2]. ∎
Let $M$ be an object in a model category. Recall that $M$ is called cofibrant if $0\rightarrow M$ is a cofibration, and it is called fibrant if $M\rightarrow 0$ is a fibration. For the Frobenius category $\mathcal{C}of$, whose objects are both cofibrant and fibrant, there is a stable category. Moreover, for the model category $\mathcal{F}ib$ there is an associated homotopy category $\mathrm{Ho}(\mathcal{F}ib)$, which is obtained by formally inverting weak equivalences, that is, the localization of $\mathcal{F}ib$ with respect to the class of weak equivalences.
The following is immediate from a fundamental result about model categories, see for example [30, Theorem 1.2.10].
@@
Note that weak equivalences are crucial in the model category. We have the following characterization, which might be of independent interest.
@@
Let ${\rm StMod}(RG)$ be the stable module category with all fibrant $RG$-modules as objects, and morphisms for any objects $M$ and $N$ given by complete cohomology of degree zero, that is,
@@
Let $M$ and $N$ be any fibrant $RG$-modules. By the completeness of the cotorsion pair $(\mathcal{C}of,\mathcal{W})$, there exists an exact sequence $0\rightarrow K_{M}\rightarrow Q(M)\rightarrow M\rightarrow 0$, where $Q(M)$ is cofibrant and $K_{M}\in\mathcal{W}$. Hence, the cofibrant approximation $Q(M)\rightarrow M$ is also a trivial fibration, and we refer to it (or simply, to $Q(M)$) as being a cofibrant replacement of $M$. Then, we have the following isomorphisms
@@
where the first one follows by [30, Theorem 1.2.10(ii)] and Proposition 6.5, and the second one holds by Lemma 6.7.
By basic properties of cofibrant modules (see Proposition 5.5), for fibrant $RG$-modules $M$ and $N$, there exists an integer $r>>0$, such that both $\Omega^{r}(M)$ and $\Omega^{r}(N)$ are cofibrant modules, and moreover, the projective dimension of $K_{M}$ and $K_{N}$ are not more than $r-1$. For $M$, we have the following commutative diagram with exact rows
[…]
where $P_{i}$ and $P^{\prime}_{i}$ are all projective $RG$-modules; this yields an exact sequence
$0\longrightarrow\Omega^{r}(M)\longrightarrow P_{r-1}\oplus P^{\prime}_{r}\longrightarrow\cdots\longrightarrow P_{0}\oplus P^{\prime}_{1}\longrightarrow Q(M)\longrightarrow 0$
Similarly, we obtain such an exact sequence for $Q(N)$. Moreover, since $P_{0}\oplus P^{\prime}_{1}$, $\cdots$, $P_{r-1}\oplus P^{\prime}_{r}$ are projective modules, we get the following commutative diagram:
@@
Let ${\rm K}_{ac}(RG\text{-}{\rm Proj})$ and ${\rm K}_{tac}(RG\text{-}{\rm Proj})$ denote the homotopy categories of acyclic complexes and totally acyclic complexes of projective $RG$-modules, respectively. There is a functor $\Omega:{\rm K}_{tac}(RG\text{-}{\rm Proj})\rightarrow\underline{\mathcal{GP}}(RG)$ given by taking the kernel of the boundary map in degree zero.
We summarize some equivalences of triangulated categories as follows, which generalize [33, Theorem 3.10] where the equivalences
@@
Acknowledgements. The author is grateful to Professor I. Emmanouil for sharing the preprint [24] and for helpful comments and suggestions for an early version of the manuscript, and to Professors X.-W. Chen, D.-S. Li and W.-H. Qian for their helpful suggestions. The author is also grateful to the referee for helpful suggestions which result in a significant improvement of the paper. This work was supported by the National Natural Science Foundation of China (No. 11871125).